OpenStudy (anonymous):

An equation and the steps to solve it are shown below: 3x 1/2 (8x - 2) = 4 Step 1: 3x - 4x + 1 = 4 Step 2: x + 1 = 4 Step 3: x + 1 - 1 = 4 - 1 Step 4: x = 3 Which step was the error made and why? (1 point) Question 19 options: 1) Step 1, because incorrect signs have been used 2) Step 2, because the terms have been added incorrectly 3) Step 3, because numbers have been added to both sides of the equation 4) No error was made

OpenStudy (anonymous):

Fan and medal

OpenStudy (jack1):

what are your thoughts...?

OpenStudy (anonymous):

1 right 2 worng

OpenStudy (jack1):

\(\huge 3x + 1/2 (8x - 2) = 4\) is this ur original equation...?

OpenStudy (jack1):

or is it \(\huge 3x \frac{(8x - 2)}2 = 4\)

OpenStudy (jack1):

anyway i think step 1, as incorrecet signs (the -2 changed to a +1)

OpenStudy (anonymous):

ops then I got it wrong on my quiz i just entered it

OpenStudy (anonymous):

2 people say its 1 wrong and 1 person 2 is wrong. I will agree with you